What is the slope of the line 2x + 3y + 3 = 0?
Answer by prateek(6)

the form of the the equation is Ax+By+C=0
the slope the line is =-A/B

where A =Coefficient of the x,
B =Coefficient of the y
equation is the 2x+3y+3=0
hence, the slope of the line is ,

m=-2/3